# Imperative Learning: A Self-supervised Neuro-Symbolic Learning Framework for Robot Autonomy

### v6

@@ 2.3 Neuro-symbolic Learning
aims to inject interpretable and deterministic logical rules into neural networks (Serafini and Garcez, 2016; Riegel et al., 2020; Xie et al., 2019; Ignatiev et al., 2018). Some previous work directly obtained such knowledge from human expert (Xu et al., 2018; Xie et al., 2019, 2021; Manhaeve et al., 2018; Riegel et al., 2020; Yang et al., 2020) or an oracle for controllable deduction in neural networks (Mao et al., 2018; Wang et al., 2022; Hsu et al., 2023), refered as deductive methods. Representative works include DeepProbLog (Manhaeve et al., 2018), logical neural network (Riegel et al., 2020), and semantic Loss (Xu et al., 2018). Despite their success, deductive methods require structured formal symbolic knowledge from humans, which is not always available. Besides, they still suffer from scaling to larger and more complex problems. In contrast, inductive approaches induct the structured symbolic representation for semi-supervised efficient network learning. One popular strategy is based on forward-searching algorithms (Li et al., 2020b, 2022b; Evans and Grefenstette, 2018; Sen et al., 2022), which is time-consuming and hard to scale up. Others borrow gradient-based neural networks to conduct rule induction, such as SATNet (Wang et al., 2019), NeuralLP (Yang et al., 2017), and neural logic machine (NLM) (Dong et al., 2019). In particular, NLM introduced a new network architecture inspired by first-order logic, which shows better compositional generalization capability than vanilla neural nets. However, existing inductive algorithms either work on structured data like knowledge graphs (Yang et al., 2017; Yang and Song, 2019) or only experiment with toy image datasets (Shindo et al., 2023; Wang et al., 2019). We push the limit of this research into robotics with IL, providing real-world applications with high-dimensional image data.
@@
in NeSy learning refers to approaches that integrate physical knowledge into the learning process to improve interpretability, generalization, and safety. Two major paradigms in this space are physics-informed and physics-infused learning. Physics-informed learning typically incorporates physical laws into the training objective by adding physics-based loss terms or constraints, guiding neural networks to produce physically consistent outputs (Raissi et al., 2019; Karniadakis et al., 2021). For example, in fluid dynamics, the Navier–Stokes equations can be enforced as soft penalties to regularize flow predictions (Sun and Wang, 2020), and in structural mechanics, variational formulations are used to encode stress-strain behavior (Rao et al., 2021). In contrast, physics-infused learning embeds physical laws or physics-based models directly into the model architecture (Vargas Venegas and Huang, 2023; Martin-Linares et al., 2023). For instance, Romero et al. (2023); Han et al. (2024) incorporated differentiable trajectory optimization into learning-based planning and control pipelines, allowing end-to-end learning with explicit physical reasoning; Zhao et al. (2024) programmed energy conservation laws into the model architecture for motion prediction in off-road driving. These methods tightly couple data-driven components with physical structure, enabling more expressive and interpretable models that can propagate gradients through physical interactions. We aim to incorporate physical reasoning via differentiable optimization within a unified neuro-symbolic learning framework for robot autonomy.
@@ 3.2 Formulation
One of the most important properties of the framework in Figure 1 is that the neural, reasoning, and memory modules can perform reciprocal learning in a self-supervised manner. This is achieved by formulating this framework as a BLO. Denote the neural system as $\bm{z}=f({\bm{\theta}},\bm{x})$, where $\bm{x}$ represents the sensor measurements, ${\bm{\theta}}$ represents the perception-related learnable parameters, and $\bm{z}$ represents the neural outputs such as semantic attributes; the reasoning engine as $g(f,M,{\bm{\mu}})$ with reasoning-related parameters ${\bm{\mu}}$ and the memory system as $M({\bm{\gamma}},{\bm{\nu}})$, where ${\bm{\gamma}}$ is perception-related memory parameters (Wang et al., 2021a) and ${\bm{\nu}}$ is reasoning-related memory parameters. In this context, our imperative learning (IL) is formulated as a special BLO:
@@
[…]
where $\xi$ is a general constraint (either equality or inequality); $U$ and $L$ are the upper-level (UL) and lower-level (LL) cost functions; and $\bm{\psi}\doteq[{\bm{\theta}}^{\top},{\bm{\gamma}}^{\top}]^{\top}$ are stacked UL variables and $\bm{\phi}\doteq[{\bm{\mu}}^{\top},{\bm{\nu}}^{\top}]^{\top}$ are stacked LL variables, respectively. Alternatively, $U$ and $L$ are also referred to as the neural cost and symbolic cost, respectively. As aforementioned, the term “imperative” is used to denote the passive nature of the learning process: once optimized, the neural system $f$ in the UL cost will be driven to align with the LL reasoning engine $g$ (e.g., logical, physical, or geometrical reasoning process) with constraint $\xi$, so that it can learn to generate logically, physically, or geometrically feasible semantic attributes or predicates. In some applications, $\bm{\psi}$ and $\bm{\phi}$ are also referred to as neuron-like and symbol-like parameters, respectively.
@@
As presented in Section 1, the formulation of IL is motivated by one important observation: many symbolic reasoning engines including geometric, physical, and logical reasoning, can be optimized or solved without providing labels. This is evident in methods like logic reasoning like equation discovery (Billard and Diday, 2002) and A search (Hart et al., 1968), geometrical reasoning such as bundle adjustment (BA) (Agarwal et al., 2010), and physical reasoning like model predictive control (Kouvaritakis and Cannon, 2016). The IL framework leverages this phenomenon and jointly optimizes the three modules by BLO, which enforces the three modules to mutually correct each other. Consequently, all three modules can learn and evolve in a self-supervised manner by observing the world. However, it is worth noting that, although IL is designed for self-supervised learning, it can easily adapt to supervised or weakly supervised learning by involving labels either in UL or LL cost functions or both.
@@
The memory system within the IL framework is a general component that can retain and retrieve information online. Specifically, it can be any structure that is associated with write and read operations to retain and retrieve data (Wang et al., 2021a). A memory can be a neural network, where information is “written” into the parameters and is “read” through a set of math operations or implicit mapping, e.g., a neural radiance fields (NeRF) model (Mildenhall et al., 2021); It can also be a structure with explicit physical meanings such as a map created online, a set of logical rules inducted online, or even a dataset collected online; It can also be the memory system of LLMs in textual form, such as retrieval-augmented generation (RAG) (Lewis et al., 2020), which writes, reads, and manages symbolic knowledge. Hyperdimensional computing (Kleyko et al., 2022, 2023) also presents a promising direction for memory modeling, though it falls outside the scope of this work. In this article, we will explore several customized memory structures tailored to specific applications, as detailed in Section 4.
@@ 3.3 Optimization
[…]
1:  while Not Convergent do
3:     Efficient estimation of UL gradients in (2) via Unrolled Differentiation: Compute ^$\hat{\nabla}_{\bm{\theta}}U$ and ^$\hat{\nabla}_{\bm{\gamma}}U$ by having $\bm{\Phi}$ in (7) and ${U}(\bm{\psi},\bm{\Phi}(\bm{\psi}))$ in (8) via AutoDiff.Implicit Differentiation (Algorithm 2): Compute
^$\hat{\nabla}_{\bm{\psi}}U=\frac{\partial U}{\partial\bm{\psi}}\Big{|}_{\bm{\phi}_{T}}+\frac{\partial U}{\partial\bm{\phi}^{*}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\partial\bm{\phi}^{*}}{\partial{\bm{\psi}}}}\Big{|}_{\bm{\phi}_{T}},$
4:  end while
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\partial\bm{\phi}^{*}}{\partial\bm{\psi}}}=\left[\begin{aligned} \frac{\partial\bm{\mu}^{*}}{\partial\bm{\theta}}&\quad\frac{\partial\bm{\mu}^{*}}{\partial\bm{\gamma}}\\ \frac{\partial\bm{\nu}^{*}}{\partial\bm{\theta}}&\quad\frac{\partial\bm{\nu}^{*}}{\partial\bm{\gamma}}\\ \end{aligned}\right],\$ (3)
@@
where $\frac{\partial U}{\partial\bm{\psi}}=\left[(\frac{\partial U}{\partial f}\frac{\partial f}{\partial{\bm{\theta}}})^{\top},(\frac{\partial U}{\partial M}\frac{\partial M}{\partial{\bm{\gamma}}})^{\top}\right]^{\top}$ and $\frac{\partial U}{\partial\bm{\phi}^{*}}=\left[(\frac{\partial U}{\partial g}\frac{\partial g}{\partial{\bm{\mu}}^{*}})^{\top},(\frac{\partial U}{\partial M}\frac{\partial M}{\partial{\bm{\nu}}^{*}})^{\top}\right]^{\top}$. There are typically two methods for calculating those gradients, i.e., unrolled differentiation and implicit differentiation. We summarize a generic framework incorporating both methods in Algorithm 1 to provide a clearer understanding.
@@ 3.3.1 Unrolled Differentiation
where $\frac{\partial\bm{\Phi}(\bm{\psi})}{\partial\bm{\psi}}$ can be computed via AutoDiff instead of directly calculating the four terms in $\frac{\partial\bm{\phi}^{*}}{\partial\bm{\psi}}$ of (3).
@@
3.3.2 Implicit Differentiation
^$\hat{\nabla}_{\bm{\psi}}U\approx\frac{\partial U}{\partial\bm{\psi}}\Big{|}_{\bm{\phi}_{T}}+\frac{\partial U}{\partial\bm{\phi}^{*}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\partial\bm{\phi}^{*}}{\partial{\bm{\psi}}}}\Big{|}_{\bm{\phi}_{T}}.$ (11)
Solving the equation gives us the implicit gradients as
where $\bm{H}_{\bm{\phi}^{*}\bm{\phi}^{*}}$ is denoted as $\bm{H}$ for simplicity. The linear system (15) can be solved without explicitly calculating and storing the Hessian matrix by solvers such as conjugate gradient (Hestenes and Stiefel, 1952) or gradient descent. For example, due to the gradient $\frac{\partial Q(\bm{q})}{\partial\bm{q}}=\bm{H}\bm{q}-\bm{v}$, the updating scheme based on the gradient descent algorithm is:
$\bm{q}_{k}=\bm{q}_{k-1}-\eta\left(\bm{H}\bm{q}_{k-1}-\bm{v}\right).$ (16)
@@ 3.3.2 Implicit Differentiation
Implicit differentiation is complicated to implement but there is one approximation, which is to ignore the implicit components and only use the direct part ^$\hat{\nabla}_{\bm{\psi}}U\approx\frac{\partial U}{\partial\bm{\psi}}\Big{|}_{\bm{\phi}_{T}}$. This is equivalent to taking the solution ${\bm{\phi}}_{T}$ from the LL optimization as constants in the UL problem. Such an approximation is more efficient but introduces an error term
@@
Second, perception-related (neuron-like) parameters are usually updated using gradient descent algorithms such as SGD (Sutskever et al., 2013). However, such simple first-order optimization methods are often inadequate for LL symbolic reasoning, e.g., geometric problems (Wang et al., 2023) usually need second-order optimizers. Therefore, separating neuron-like parameters from reasoning-related parameters makes the BLO in IL easier to solve and analyze. However, this again depends on the LL tasks.
@@
The optimality for generic non-convex LL problems remain an open challenge and are sometimes NP-hard (Dempe et al., 2015) in BLO theory. This difficulty arises because non-uniqueness in the LL problem prevents the implicit differentiation theorem from guaranteeing the existence of implicit differentiability of $\phi^{*}$ with respect to the UL variables $\psi$. A weaker condition for establishing optimality is the Polyak-Lojasiewicz (PL) condition (Shen and Chen, 2023), which allows for an overall non-convex landscape while maintaining certain convexity-like structures. Such structures have been observed in overparameterized neural networks and can guarantee an $\epsilon$-approximate solution to the BLO problem (Liu et al., 2020a). We will show more detailed theoretical guarantees and their assumptions in the next section with respect to the types of LL optimization. To ensure the language remaining accessible and not overly theoretical, we will only provide summarized results and provide references to their sources. In this way, we hope it will be helpful for readers to easily find more relevant formal results when needed.

### v7

@@ 2.3 Neuro-symbolic Learning
aims to inject interpretable and deterministic logical rules into neural networks (serafini2016logic; riegel2020logical; xie2019embedding; ignatiev2018pysat). Some previous work directly obtained such knowledge from human expert (xu2018semantic; xie2019embedding; xie2021embedding; manhaeve2018deepproblog; riegel2020logical; ijcai2020p243) or an oracle for controllable deduction in neural networks (mao2018neuro; wang2022programmatically; hsu2023ns3d), refered as deductive methods. Representative works include DeepProbLog (manhaeve2018deepproblog), logical neural network (riegel2020logical), and semantic Loss (xu2018semantic). Despite their success, deductive methods require structured formal symbolic knowledge from humans, which is not always available. Besides, they still suffer from scaling to larger and more complex problems. In contrast, inductive approaches induct the structured symbolic representation for semi-supervised efficient network learning. One popular strategy is based on forward-searching algorithms (li2020closed; li2022softened; evans2018dILP; sen2022loa), which is time-consuming and hard to scale up. Others borrow gradient-based neural networks to conduct rule induction, such as SATNet (wang2019satnet), NeuralLP (yang2017neurallp), and neural logic machine (NLM) (dong2019nlm). In particular, NLM introduced a new network architecture inspired by first-order logic, which shows better compositional generalization capability than vanilla neural nets. However, existing inductive algorithms either work on structured data like knowledge graphs (yang2017neurallp; yang2019nlil) or only experiment with toy image datasets (shindo2023aILP; wang2019satnet). We push the limit of this research into robotics with IL, providing real-world applications with high-dimensional image data.
@@
in NeSy learning refers to approaches that integrate physical knowledge into the learning process to improve interpretability, generalization, and safety. Two major paradigms in this space are physics-informed and physics-infused learning. Physics-informed learning typically incorporates physical laws into the training objective by adding physics-based loss terms or constraints, guiding neural networks to produce physically consistent outputs (raissi2019physics; karniadakis2021physics). For example, in fluid dynamics, the Navier–Stokes equations can be enforced as soft penalties to regularize flow predictions (sun2020physics), and in structural mechanics, variational formulations are used to encode stress-strain behavior (rao2021physics). In contrast, physics-infused learning embeds physical laws or physics-based models directly into the model architecture (vargas2023physics; martin2023physics). For instance, romero2023actor; han2024learning incorporated differentiable trajectory optimization into learning-based planning and control pipelines, allowing end-to-end learning with explicit physical reasoning; zhao2024physord programmed energy conservation laws into the model architecture for motion prediction in off-road driving. These methods tightly couple data-driven components with physical structure, enabling more expressive and interpretable models that can propagate gradients through physical interactions. We aim to incorporate physical reasoning via differentiable optimization within a unified neuro-symbolic learning framework for robot autonomy.
@@ 3.2 Formulation
One of the most important properties of the framework in Figure 1 is that the neural, reasoning, and memory modules can perform reciprocal learning in a self-supervised manner. This is achieved by formulating this framework as a BLO. Denote the neural system as $\bm{z}=f({\bm{\theta}},\bm{x})$, where $\bm{x}$ represents the sensor measurements, $\bm{\theta}$} represents the perception-related learnable parameters, and $\bm{z}$ represents the neural outputs such as semantic attributes; the reasoning engine as $g(f,M,{\bm{\mu}})$ with reasoning-related parameters $\bm{\mu}$} and the memory system as $M({\bm{\gamma}},{\bm{\nu}})$, where $\bm{\gamma}$} is perception-related memory parameters (wang2021unsupervised) and $\bm{\nu}$} is reasoning-related memory parameters. In this context, our imperative learning (IL) is formulated as a special BLO:
@@
$\textrm{s.t.}\quad\xi(M({\bm{\gamma}},{\bm{\nu}}),{\bm{\mu}},f({\bm{\theta}},\bm{x}))=\text{ or }\leq 0,$ (1c)
where $\xi$ is a general constraint (either equality or inequality); $U$ and LL are the upper-level (UL) and lower-level (LL) cost functions; and $\bm{\psi}\doteq[{\bm{\theta}}^{\top},{\bm{\gamma}}^{\top}]^{\top}$ are stacked UL variables and $\bm{\phi}\doteq[{\bm{\mu}}^{\top},{\bm{\nu}}^{\top}]^{\top}$ are stacked LL variables, respectively. Alternatively, $U$ and LL are also referred to as the neural cost and symbolic cost, respectively. As aforementioned, the term “imperative” is used to denote the passive nature of the learning process: once optimized, the neural system $f$ in the UL cost will be driven to align with the LL reasoning engine $g$ (e.g., logical, physical, or geometrical reasoning process) with constraint $\xi$, so that it can learn to generate logically, physically, or geometrically feasible semantic attributes or predicates. In some applications, $\bm{\psi}$ and $\bm{\phi}$ are also referred to as neuron-like and symbol-like parameters, respectively.
@@
As presented in Section 1, the formulation of IL is motivated by one important observation: many symbolic reasoning engines including geometric, physical, and logical reasoning, can be optimized or solved without providing labels. This is evident in methods like logic reasoning like equation discovery (billard2002symbolic) and A search (hart1968formal), geometrical reasoning such as bundle adjustment (BA) (agarwal2010bundle), and physical reasoning like model predictive control (kouvaritakis2016model). The IL framework leverages this phenomenon and jointly optimizes the three modules by BLO, which enforces the three modules to mutually correct each other. Consequently, all three modules can learn and evolve in a self-supervised manner by observing the world. However, it is worth noting that, although IL is designed for self-supervised learning, it can easily adapt to supervised or weakly supervised learning by involving labels either in UL or LL cost functions or both.
@@
The memory system within the IL framework is a general component that can retain and retrieve information online. Specifically, it can be any structure that is associated with write and read operations to retain and retrieve data (wang2021unsupervised). A memory can be a neural network, where information is “written” into the parameters and is “read” through a set of math operations or implicit mapping, e.g., a neural radiance fields (NeRF) model (mildenhall2021nerf); It can also be a structure with explicit physical meanings such as a map created online, a set of logical rules inducted online, or even a dataset collected online; It can also be the memory system of LLMs in textual form, such as retrieval-augmented generation (RAG) (lewis2020retrieval), which writes, reads, and manages symbolic knowledge. Hyperdimensional computing (kleyko2022survey; kleyko2023survey) also presents a promising direction for memory modeling, though it falls outside the scope of this work. In this article, we will explore several customized memory structures tailored to specific applications, as detailed in Section 4.
@@ 3.3 Optimization
BLO has been explored in frameworks such as meta-learning (finn2017model), hyperparameter optimization (franceschi2018bilevel), and reinforcement learning (hong2020two). However, most of the theoretical analyses have primarily focused on their applicability to data-driven models, where 1st-order gradient descent (GD) is frequently employed (ji2021bilevel; gould2021deep). Nevertheless, many reasoning tasks present unique challenges that make GD less effective. For example, geometrical reasoning like BA requires 2nd-order optimizers (fu2024islam) such as Levenberg-Marquardt (LM) (marquardt1963algorithm); multi-robot routing needs combinatorial optimization over discrete variables (ren2023cbss). Employing such LL optimizations within the BLO framework introduces extreme complexities and challenges, which are still underexplored (ji2021bilevel). Therefore, we will first delve into general BLO and then provide practical examples covering distinct challenges of LL optimizations in our IL framework.
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\partial\bm{\phi}^{*}}{\partial\bm{\psi}}}=\left[\begin{aligned} \frac{\partial\bm{\mu}^{*}}{\partial\bm{\theta}}&\quad\frac{\partial\bm{\mu}^{*}}{\partial\bm{\gamma}}\\ \frac{\partial\bm{\nu}^{*}}{\partial\bm{\theta}}&\quad\frac{\partial\bm{\nu}^{*}}{\partial\bm{\gamma}}\\ \end{aligned}\right],\\$ (3)
@@
where $\frac{\partial U}{\partial\bm{\psi}}=\left[(\frac{\partial U}{\partial f}\frac{\partial f}{\partial{\bm{\theta}}})^{\top},(\frac{\partial U}{\partial M}\frac{\partial M}{\partial{\bm{\gamma}}})^{\top}\right]^{\top}$ and $\frac{\partial U}{\partial\bm{\phi}^{*}}=\left[(\frac{\partial U}{\partial g}\frac{\partial g}{\partial{\bm{\mu}}^{*}})^{\top},(\frac{\partial U}{\partial M}\frac{\partial M}{\partial{\bm{\nu}}^{*}})^{\top}\right]^{\top}$. Computing these gradients is challenging. We therefore first introduce a simple yet effective method, alternating optimization, which avoids explicit gradient computation, and then present unrolled differentiation and implicit differentiation to compute them.
3.3.1 Alternating Optimization
One simple strategy for BLO is alternating optimization, which avoids explicit computation of the implicit gradients $\frac{\partial\bm{\phi}^{*}}{\partial\bm{\psi}}$. Specifically, the UL and LL problems are solved in an alternating and iterative manner. Given UL variables $\bm{\psi}^{(k)}$, the LL problem is first (approximately) solved to obtain $\bm{\phi}^{(k)}$, possibly under constraints (1c). Then, fixing $\bm{\phi}^{(k)}$, the UL variables are updated by minimizing the UL cost $U$ using any standard gradient-based methods. This process is repeated iteratively until convergence, as illustrated in (5).
[…]
Alternating optimization can be viewed as a block coordinate descent on the bilevel structure. The LL reasoning engine acts as a constraint-enforcing module that corrects neural and memory outputs, while the UL learning process gradually aligns these modules with the LL solutions. A key advantage of alternating optimization is that it places minimal requirements on the LL solver. The LL problem need not be differentiable and may involve discrete variables or non-smooth constraints. This makes alternating optimization particularly suitable for symbolic reasoning tasks that rely on combinatorial search, geometric optimization, physics-based solvers, or other settings where implicit differentiation is difficult to compute. Alternating optimization has been verified to be effective in practice, such as in geometric perception (yang2021self) and rule induction in Section 5.2. However, since this approach ignores the dependency of $\bm{\phi}^{*}$ on $\bm{\psi}$ and thus yields biased gradients, it can sometimes be unstable and sensitive to initialization.
@@ 3.3.2 Unrolled Differentiation
where $\frac{\partial\bm{\Phi}(\bm{\psi})}{\partial\bm{\psi}}$ can be computed via AutoDiff instead of directly calculating the four terms in $\frac{\partial\bm{\phi}^{*}}{\partial\bm{\psi}}$ of (3). A generic framework incorporating both unrolled differentiation and implicit differentiation is listed in Algorithm 1 for clarity.
1:  while Not Convergent do
3:   Efficient estimation of UL gradients in (2) via Unrolled Differentiation: Compute $\hat{\nabla}_{\bm{\theta}}U$ and $\hat{\nabla}_{\bm{\gamma}}U$ by having $\bm{\Phi}$ in (8) and ${U}(\bm{\psi},\bm{\Phi}(\bm{\psi}))$ in (9) via AutoDiff.Implicit Differentiation (Algorithm 2): Compute
$\hat{\nabla}_{\bm{\psi}}U=\frac{\partial U}{\partial\bm{\psi}}\Big|_{\bm{\phi}_{T}}+\frac{\partial U}{\partial\bm{\phi}^{*}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\partial\bm{\phi}^{*}}{\partial{\bm{\psi}}}}\Big|_{\bm{\phi}_{T}},$
4:  end while
3.3.3 Implicit Differentiation
$\hat{\nabla}_{\bm{\psi}}U\approx\frac{\partial U}{\partial\bm{\psi}}\Big|_{\bm{\phi}_{T}}+\frac{\partial U}{\partial\bm{\phi}^{*}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\partial\bm{\phi}^{*}}{\partial{\bm{\psi}}}}\Big|_{\bm{\phi}_{T}}.$ (10)
Solving the equation gives us the implicit gradients as
where $\bm{H}_{\bm{\phi}^{*}\bm{\phi}^{*}}$ is denoted as $\bm{H}$ for simplicity. The linear system (14) can be solved without explicitly calculating and storing the Hessian matrix by solvers such as conjugate gradient (hestenes1952methods) or gradient descent. For example, due to the gradient $\frac{\partial Q(\bm{q})}{\partial\bm{q}}=\bm{H}\bm{q}-\bm{v}$, the updating scheme based on the gradient descent algorithm is:
$\bm{q}_{k}=\bm{q}_{k-1}-\eta\left(\bm{H}\bm{q}_{k-1}-\bm{v}\right).$ (15)
@@ 3.3.3 Implicit Differentiation
Implicit differentiation is complicated to implement but there is one approximation, which is to ignore the implicit components and only use the direct part $\hat{\nabla}_{\bm{\psi}}U\approx\frac{\partial U}{\partial\bm{\psi}}\Big|_{\bm{\phi}_{T}}$. This is equivalent to taking the solution ${\bm{\phi}}_{T}$ from the LL optimization as constants in the UL problem. Such an approximation is more efficient but introduces an error term
@@
Second, perception-related (neuron-like) parameters are usually updated using gradient descent algorithms such as SGD (sutskever2013importance). However, such simple first-order optimization methods are often inadequate for LL symbolic reasoning, e.g., geometric problems (wang2023pypose) usually need second-order optimizers. Therefore, separating neuron-like parameters from reasoning-related parameters makes the BLO in IL easier to solve and analyze. However, this again depends on the LL tasks.
@@
The optimality for generic non-convex LL problems remain an open challenge and are sometimes NP-hard (dempe2015bilevel) in BLO theory. This difficulty arises because non-uniqueness in the LL problem prevents the implicit differentiation theorem from guaranteeing the existence of implicit differentiability of $\phi^{*}$ with respect to the UL variables $\psi$. A weaker condition for establishing optimality is the Polyak-Lojasiewicz (PL) condition (shen2023penalty), which allows for an overall non-convex landscape while maintaining certain convexity-like structures. Such structures have been observed in overparameterized neural networks and can guarantee an $\epsilon$-approximate solution to the BLO problem (liu2020toward). We will show more detailed theoretical guarantees and their assumptions in the next section with respect to the types of LL optimization. To ensure the language remaining accessible and not overly theoretical, we will only provide summarized results and provide references to their sources. In this way, we hope it will be helpful for readers to easily find more relevant formal results when needed.